Find a confidence interval for μ assuming that each sample is from a normal population. (Round the value of t to 3 decimal places and your final answers to 2 decimal places.) (a) x⎯⎯ x ¯ = 25, s = 5, n = 7, 90 percent confidence. The 90% confidence interval is to (b) x⎯⎯ x ¯ = 50, s = 4, n = 19, 99 percent confidence. The 99% confidence interval is to (c) x⎯⎯ x ¯ = 121, s = 14, n = 29, 95 percent confidence. The 95% confidence interval is to
Answer #1

We know that, the confidence interval for population mean when is unknown given by, х + s tn-1,0% = 25, S = 5, n=7, 1-9 = 0.9© x=121, 5=14, n=29, 1-x=0.95 2x=0.05 .: The 95% C.I is, $ 121 14 x 2.048 129 Here, to-1,*/ t 2gt, ovos t28, 2 28, 0.025 = -